OpenStudy (anonymous):

Write a line parallel to y=4x-7 and goes through point (4,7)

OpenStudy (anonymous):

parallel line m=4 y-y1=m(x-x1) y-7=4(x-4) y-7=4x-16 y=4x-9

OpenStudy (anonymous):

Hey what about write a line perpendicular to y=4x-7 and goes through (0,-7)

OpenStudy (anonymous):

perpendicular line m= -1/4 put the x and y coordinates like I just did and input m to -1/4 solve it and bla bla